OpenStudy (anonymous):

Can someone show me how to expand log (x^3sqrt(x^2+1))/(y^2(x+1)^4)

OpenStudy (anonymous):

log (x^3sqrt(x^2+1))/(y^2(x+1)^4) do you think the solution is like this? (x^3sqrt(x^2+1) =log ---------------------- ..........................yes or no? (y^2(x+1)^4)) =log (x^3sqrt(x^2+1) - log (y^2(x+1)^4)) ................yes or no? =log (x^3) + log(x^2+1)^1/2 - [ log (y^2) + log(x+1)^4))] ...... yes or no? =3log (x) +(1/2) log(x^2+1) - [2 log (y) + 4log(x+1)].......... yes or no? =3log (x) +(1/2) log(x^2+1) - 2 log (y) - 4log(x+1)..............final answer,,,,.yes or no?

OpenStudy (anonymous):

I honestly don't know how to even do it.. the equation looks like this: \[Expand:\log\frac{ x^3\sqrt{x^2+1} }{ y^2(x+1)^4 }\]

OpenStudy (anonymous):

yup the problem i copied is correct based on what you wrote now

OpenStudy (anonymous):

okay. so the final answer looks like this? : \[3\log(x)+\frac{ 1 }{ 2 }\log(x^2+1)-2\log(y)-4\log(x+1)\]

OpenStudy (anonymous):

yes you are correct.... :D

OpenStudy (anonymous):

Thanks so much! Do you by any chance know how to write a ln equation into a single ln? ie : 2ln(x+1)-ln(x+2)-3lnx

OpenStudy (anonymous):

ypu do the same process backward its the same thing

OpenStudy (anonymous):

2ln(x+1)-ln(x+2)-3lnx check if the solution is correct =ln(x+1)^2 - ln(x+2)-lnx^3 separate all the negative signs to the right =ln(x+1)^2 -( ln(x+2)+lnx^3) ....... yes or no? l(x+1)^2 = ln ------------- ...................... yes or no? (x+2)(x^3)

OpenStudy (anonymous):

why is there an l infront of (x+1)^2 for the final answer?

OpenStudy (anonymous):

@mark_o.

OpenStudy (anonymous):

ln means natural logarithm of base e while log means log to the base 10

OpenStudy (anonymous):

ohhhh. okay i get it thank you so much for your help!!

OpenStudy (anonymous):

ok yw... have fun.... good luck now :D